The lengths of pregnancies in a small rural village are normally distributed with a mean of 270 days and a standard deviation of 15 days. A distribution of values is normal with a mean of 270 and a standard deviation of 15.
What percentage of pregnancies last fewer than 257 days?
P(X < 257 days) = %
Enter your answer as a percent accurate to 1 decimal place (do not enter the "%" sign). Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.

Answer by Theo(13342)

population mean is 270 days.
population standard deviation is 15 days.
test value = 257 days.

use z-score formula.
z = (x - m) / 15 becomes z = (257 - 270) / 15 = -.86667
z is the z-score
x is the test value
m is the mean
s is the standard deviation.

area to the left of z-score of -.86667 = .1931.
this means there is a 19.3% probability of a pregnancy lasting less than 257 days.
you would enter 19.3 as your solution is what i think based on the instructions.